OpenStudy (anonymous):

Determine which ordered pair in the form (a, b) is a solution of the given equation. 4a – 5b = 24 (6, 0) (–5, –2) (0, 6) (–2, –5)

OpenStudy (texaschic101):

just a matter of subbing I will do the first answer choice.. 4a - 5b = 24 (6,0)....a = 6 and b = 0 now we sub 4(6) - 5(0) = 24 24 - 0 = 24 24 = 24 (correct).....this is a solution....The solution would have to be the first one I tried..lol